OpenStudy (anonymous):

5 times 6=5 times 4 pluse 5 times b=?

OpenStudy (anonymous):

5*6 = 5*4 + 5*b 30 = 20 + 5b 5b = 30 - 20 5b = 10 b = 10/5 b = 2